The Length of a rectangular garden is 7 feet longer than its width. The perimeter of the garden is 38 feet. Find the width and length of the garden.
Answer by Paul(988)

Width = x
Length = x+7
Formula:
2(l+w)=P

2(x+x+7)=38
4x+14=38
4x=24
x=6

6+7=13

Hence, the length is 13ft long and the width is 6ft long.
Paul.